OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + 9x = 8 to the nearest hundredth?

OpenStudy (amorfide):

I would recommend getting it into the form ax^2+bx+c=0 then use the quadratic formula can you do that?

OpenStudy (anonymous):

not at all x.x

OpenStudy (amorfide):

okay to get into the form ax^2+bx+c=0 you must subtract 8 to both sides. this gives 2x^2 + 9x -8=0 quadratic formula, do you know what it is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x ≈ 1.22 and x ≈ 3.28 x ≈ −0.76 and x ≈ 5.26 x ≈ −1.22 and x ≈ −3.28 x ≈ −5.26 and x ≈ 0.76 These are the answers I am given that is why I do no understand

OpenStudy (amorfide):

okay well the standard form is ax^2+bx+c=0 so match a b and c with your new equation 2x^2 + 9x -8 and substitute those values into the quadratic formula